OpenStudy (anonymous):

find an equation for the inverse of the relation y=5x+3

OpenStudy (anonymous):

like solving \[5x+3=12\] subtract 3 to get \[5x=9\] then divide by 5 to get \[x=\frac{9}{5}\] in this case switch x and y and solve \[5y+3=x\] solve for y and get \[5y=x-3\] \[y=\frac{x-3}{5}\]

OpenStudy (anonymous):

thank you!